DFT TEST STUDY ON VAN DER WAALS DIMERS

\underline{ANTONIO J. HERNANDEZ}, Department of Chemistry, Simon Bolivar University, Caracas 1080A, Venezuela; CARLOS MANZANARES I., Department of Chemistry and Biochemistry, Baylor University, Waco, TX 76798; and MARY C. SALAZAR, Department of Chemistry, Simon Bolivar University, Caracas 1080A, Venezuela.

Potential energy surfaces are obtained in the framework of the supermolecule approach by using density functional theory and basis sets optimized in generalized gradient approximation exchange-correlation calculations supplemented by bond functions placed at the midpoint of the van der Waals bond. They are compared to surfaces obtained using single and double excitation coupled-cluster theory with noniterative treatment of triple excitations CCSD(T) and correlationconsistent basis sets supplemented by bond functions. Bonding properties and Vertical excitation electronic spectra of N_{2} − He and CO^{+} − He van der Waals molecules will be presented and compared to recent theoretical results.